OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x + y = 10 y = x + 8

OpenStudy (anonymous):

need help please

OpenStudy (anonymous):

x+x+8=10 2x=-2 x=-1 substitute for y

OpenStudy (llort):

(1, 9)

OpenStudy (anonymous):

y=-1+8 y=7 (-1,7)